OpenStudy (anonymous):

The number of years, N(r), since two independently evolving languages split off from a common ancestral language is approximated by N(r)=-5000 In (r), where r is the percent of words from the ancestral language common to both languages. Find N if r=80%

OpenStudy (anonymous):

I know how to set it up. It should look like N(r)=-5000 In (.80) Though I am not sure what In stands for because my homework doesn't tell me. There is an example that my homework gave me if you would like me to write it down really fast and I can't figure out where they got their answer.

OpenStudy (perl):

ln stands for the natural logarithm

OpenStudy (anonymous):

oh! what is the natural log?

OpenStudy (perl):

ln(1) means the same as \(\large \log_e(1) \)

OpenStudy (perl):

It is the logarithm with a base of a special constant \(\large e = 2.71818...\)

OpenStudy (anonymous):

\[ e^{\ln x} = x \\ \ln(e^x) = x\\ \ln(x) = \log_e(x) \]

OpenStudy (anonymous):

Though \(e^{\ln(x)}\) requires that \(x>0\).

OpenStudy (anonymous):

oh ok! so it should be N(r)=-5000(2.718181)(.80)

OpenStudy (perl):

\( \Large N(r)=-5000 \log_e (r)\) That decimal e=2.71828... never ends so its easier to just label it as e

OpenStudy (anonymous):

how do I type that into the calculator? I have a TI-84 C

OpenStudy (perl):

since r is a percent you want r / 100 80% = .80 \( \Large N(.80)=-5000 \log_e (.80)\) There should be an LN button on your calculator

OpenStudy (perl):

-5000 ln(.80)

OpenStudy (anonymous):

I know you press 2nd then the LN button but I don't know what to type in there after I press those buttons

OpenStudy (perl):

you don't need to press 2nd LN

OpenStudy (anonymous):

Oh I got it thanks!

OpenStudy (perl):

2nd LN is e^x , a different function

OpenStudy (anonymous):

so the answer should be 1115.7178 and when we round it to the nearest whole number it should be 1116 right?

OpenStudy (anonymous):

is that right?

OpenStudy (perl):

yes

OpenStudy (anonymous):

awesome, thank you for your help!